OpenStudy (anonymous):

If 10% of the group voted to change its bylaws and 10% voted against the change, how many people did not vote? The group has 200 members. PLEASE HELP!!

OpenStudy (anonymous):

Total members = 200 10% voted to change = 20 (200 * 0.10) 10% voted against the change = 20 (200 * 0.10) How many did not vote = 200 - 20 - 20 = 160

OpenStudy (anonymous):

Okay, thank you soo much that made perfect sense.